OpenStudy (anonymous):

Write a recursive formula for a sequence whose first three terms are 1, 1 and 3.

OpenStudy (anonymous):

only how to solve them

OpenStudy (anonymous):

what do you mean, how to solve them? This problem is just asking you to make up a recursive formula that fits those three numbers right?

OpenStudy (anonymous):

I was talkin to notebook sorry the question is as stands

OpenStudy (anonymous):

so its wants maybe something in the form of: \[F_{n+2}=aF_{n+1}+bF_{n}\] where a and b are integers, and it fits those three numbers.

OpenStudy (anonymous):

With your sequece (1, 1, 3), we have that: \[F_{n+2} = 3, F_{n+1} = 1, F_n = 1\] plugging this into the equation: \[3 = a(1)+b(1) \Rightarrow 3 = a + b\] So just pick numbers for a and b that sum up to three, and your good.